# M2R2: MulitModal Robotic Representation for Temporal Action Segmentation

### v1

@@ IV-A Dataset
For our experiments, we use the REASSEMBLE [13] dataset, a contact-rich manipulation dataset where a robot performs assembly and disassembly tasks with various connectors, gears, pegs, and nuts. The dataset consists of 4,551 action demonstrations across 149 recordings. It presents significant challenges for temporal action segmentation due to the high visual similarity between objects (e.g., pegs and gears that differ only slightly in size) and the high median number of actions per recording (36). We select this dataset because it includes annotations for robot temporal action segmentation and provides measurements from multiple sensors. These include multi-view RGB cameras, multiple microphones, proprioceptive robot data, and force/torque readings at the robot’s wrist. For our experiments, we use both external cameras and the gripper-mounted microphone, along with force and torque measurements, the end-effector pose, and twist. The dataset contains 69 fine-grain, each combining an activity name (e.g., ”insert”) with an object name (e.g., ”USB”). We also define a coarse labeling scheme that uses only the activity names.
@@ IV-B Metrics
The computer vision community evaluates temporal action segmentation (TAS) models using five standard metrics [11]: frame-level accuracy, EDIT score, and segmental F1 scores at 10%, 25%, and 50% overlap thresholds. Frame-level accuracy measures the percentage of correctly classified frames, while the EDIT score quantifies sequence alignment via the number of insertions, deletions, and swaps needed to match predictions to ground truth. Segmental F1 scores assess the quality of predicted action segments based on the minimum required overlap with ground truth.
Robotic TAS research employs similar metrics [4], often including the detection rate (DR) [4, 5], which evaluates action boundary detection. DR is computed as the F1 score by defining a ±$\pm t_{e}$-frame window around each ground truth boundary. A predicted boundary within this window counts as a true positive; multiple detections in the same window contribute one true positive, with the rest as false positives. Predictions outside any window are false positives, while unmatched ground truth boundaries are false negatives.
@@ IV-C Baselines
We evaluate the performance of BR-Prompt [16], a purely vision-based video feature extractor. We pretrain BR-Prompt using its default hyperparameters, sampling 16-frame windows during training. These windows typically span 2–3 actions and, at most, 6 actions. We compare our method to two state-of-the-art proprioception-based temporal action segmentation approaches used in robotics: BOCPD [4], which segments actions using the full 6DoF wrench, and AWE [1], which relies on position trajectories. We tune their hyperparameters to ensure the number of predicted segments aligns with the ground truth. Since these methods do not predict action labels, we assign each segment the ground truth class with the highest overlap. Additionally, we train three state-of-the-art deep learning models—MSTCN [8], ASRF [10], and DiffAct [11]—on M2R2 multimodal features.
@@ IV-D Quantitative Evaluation
Figure 4: Quantitative evaluation of different baseline TAS models. AWE [1] performs poor in sections of highly nonlinear movement, while BOCPD [4] tends to over-segment in areas with high variation in force interactions (marked in black). The deep learning-based approaches occasionally misidentify objects.
@@ IV-E Qualitative Evaluation
TABLE II: Modality ablation study. All models where trained with the DiffAct TAS model [11]. We evaluate on fine-grain labels.

### v2

@@ IV-A Dataset
For our experiments, we use the REASSEMBLE [sliwowski2025reassemble] dataset, a contact-rich manipulation dataset where a robot performs assembly and disassembly tasks with various connectors, gears, pegs, and nuts. The dataset consists of 4,551 action demonstrations across 149 recordings. It presents significant challenges for temporal action segmentation due to the high visual similarity between objects (e.g., pegs and gears that differ only slightly in size) and the high median number of actions per recording (36). We select this dataset because it includes annotations for robot temporal action segmentation and provides measurements from multiple sensors. These include multi-view RGB cameras, multiple microphones, proprioceptive robot data, and force/torque readings at the robot’s wrist. For our experiments, we use both external cameras and the gripper-mounted microphone, along with force and torque measurements, the end-effector pose, and twist. The dataset contains 69 fine-grain, each combining an activity name (e.g., ”insert”) with an object name (e.g., ”USB”). We also define a coarse labeling scheme that uses only the activity names.
@@ IV-B Metrics
The computer vision community evaluates temporal action segmentation (TAS) models using five standard metrics [diffact]: frame-level accuracy, EDIT score, and segmental F1 scores at 10%, 25%, and 50% overlap thresholds. Frame-level accuracy measures the percentage of correctly classified frames, while the EDIT score quantifies sequence alignment via the number of insertions, deletions, and swaps needed to match predictions to ground truth. Segmental F1 scores assess the quality of predicted action segments based on the minimum required overlap with ground truth.
Robotic TAS research employs similar metrics [BOCPD], often including the detection rate (DR) [BOCPD, EIBAND2023104367], which evaluates action boundary detection. DR is computed as the F1 score by defining a $\pm t_{e}$-frame window around each ground truth boundary. A predicted boundary within this window counts as a true positive; multiple detections in the same window contribute one true positive, with the rest as false positives. Predictions outside any window are false positives, while unmatched ground truth boundaries are false negatives.
@@ IV-C Baselines
We evaluate the performance of BR-Prompt [BRPrompt], a purely vision-based video feature extractor. We pretrain BR-Prompt using its default hyperparameters, sampling 16-frame windows during training. These windows typically span 2–3 actions and, at most, 6 actions. We compare our method to two state-of-the-art proprioception-based temporal action segmentation approaches used in robotics: BOCPD [BOCPD], which segments actions using the full 6DoF wrench, and AWE [shi2023waypointbased], which relies on position trajectories. We tune their hyperparameters to ensure the number of predicted segments aligns with the ground truth. Since these methods do not predict action labels, we assign each segment the ground truth class with the highest overlap. Additionally, we train three state-of-the-art deep learning models—MSTCN [mstcn], ASRF [asrf], and DiffAct [diffact]—on M2R2 multimodal features.
@@ IV-D Quantitative Evaluation
Figure 4: Quantitative evaluation of different baseline TAS models. AWE [shi2023waypointbased] performs poor in sections of highly nonlinear movement, while BOCPD [BOCPD] tends to over-segment in areas with high variation in force interactions (marked in black). The deep learning-based approaches occasionally misidentify objects.
@@ IV-E Qualitative Evaluation
TABLE II: Modality ablation study. All models where trained with the DiffAct TAS model [diffact]. We evaluate on fine-grain labels.